OpenStudy (anonymous):

How do you solve for log base4 x + log base4 (x+1)=1/2 ???

OpenStudy (zzr0ck3r):

\(\log_4(x) + \log_4(x+1) = \log_4(x(x+1)) = \log_4(x^2+x)\) so \(\log_4(x^2+x) = \frac{1}{2}\implies 4^{\frac{1}{2}}=x^2+x\implies x^2+x-2=0\implies x = -2,1\) but domain of log is positive numbers, so \[x=1\]

OpenStudy (anonymous):

thanks(: